OpenStudy (anonymous):

(-27/8)^(2/3) needs work to be shown

OpenStudy (jamesj):

First note that (-27/8)^(2/3) = ( (-27/8)^(1/3) )^2 So let's calculate (-27/8)^(1/3) first. (-27/8)^(1/3) = -27^(1/3) / 8 ^(1/3) = -3 / 2 because -3^3 = -27 and 2^3 = 8 Now that means that ( (-27/8)^(1/3) )^2 = ( -3 / 2)^2 = (-3)^2/2^2 = ... Can you take it from here?

OpenStudy (anonymous):

\[(\frac{-27}{8})^{\frac{2}{3}}=\frac{-3^{3^{2/3}}}{2^{3^{2/3}}}=(\frac{-3}{2})^2\]